OpenStudy (anonymous):

What is the vertex of the parabola whose equation is y = (x + 1)2 + 3? (-1, -3) (-1, 3) (1, 3) HELP

jimthompson5910 (jim_thompson5910):

Hint: the vertex of y = a(x-h)^2 + k is the point (h,k)

OpenStudy (anonymous):

help me @timo86m please

jimthompson5910 (jim_thompson5910):

for example the vertex of y = 2(x-3)^2 + 7 is (3,7) since h = 3 and k = 7

OpenStudy (anonymous):

thanks! ive got it now

jimthompson5910 (jim_thompson5910):

what did you get

OpenStudy (anonymous):

I DID THIS LAST YEAR HOLD ON

OpenStudy (anonymous):

oh its closed

OpenStudy (anonymous):

sorry i closed before you said you knew how to do it. can you tell me through message? @trevino